An average of two people per minute arrive at the check-out counter of a grocery store. What is the probability that exactly one will arrive in a given minute?
Answer #1

µ = 2 per minute

This is a poisson distribution.

P(X = x) = e * µx / x!

P(X = 1) = e-2 * 21 / 1! = 0.2707